ENVIRONMENTALLY INDUCED CHANGES IN THE BARRIERS TO INTERNAL MOTION ^{a}

\underline{TIMOTHY M. KORTER} and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

The rotationally resolved S_{1}←S_{0} fluorescence excitation spectra of N-methyl-, 3-methyl-, and 5-methylindole and their Ar van der Waals complexes have been obtained. Analyses of these spectra show that the torsional barrier for methyl group rotation is significantly influenced by the attachment of a single Ar atom to the isolated molecule. For example, N-methylindole has V_{3}(S_{0}) = 244 cm^{−1} and V_{3}(S_{1}) = 115 cm^{−1}; upon complexation, these values change to V_{3}(S_{0}) = 303 cm^{−1} and V_{3}(S_{1}) = 126 cm^{−1}. Possible reasons for this behavior will be discussed; including (a) simple steric effects,

(b) Ar atom induced changes in the electronic distribution of the parent molecule, and/or (c) changes in the nature of the motion itself.

aWork supported by NSF.